### v1

@@ 2 Related Work
We first survey the existing labeled datasets for information extraction tasks from PDFs. The earliest datasets for Document Image Analysis and Recognition (DIAR) can be traced back to the 1990s, including NIST[14] and UW[15] datasets. Over the past 30 years, there has been significant progress in this direction. Notable, since 2015, the field has seen a boom in terms of dataset availability. Initially, the datasets were small and comprised scanned PDFs or images, but as time progressed, datasets with digital-born PDFs became more prevalent. In creating a Document Layout Analysis (DLA) dataset, annotation is the most challenging part. The process of annotation can be broadly classified into three categories[16]:
@@
Table 2: Summary of studies comparing PDF parsers Evaluation metrics used by the studies: Precision(P), Recall(R), and F1 Score.
@@ 4.1 For Text Extraction:
Scientific and Patent categories presented notable challenges. In the Scientific category, all tools showed a marked decrease in performance, with Pypdfium maintaining a slight edge at an F1 score of 0.8525 and BLEU score of 0.7089 as shown in Table 4. The Patent category exhibited the widest performance gap among tools. PyMuPDF and pypdfium significantly outperformed others, scoring F1 scores of 0.973 and 0.969 respectively. The significant drop in scores for Scientific and Patent documents highlights a persistent challenge in PDF parsing technology. For Scientific documents, we compare the rule-based parser with Nougat a Visual transformer model and it outperforms all rule-based parsers by a huge margin as shown in Figure 8.
@@ 4.2 For Table Extraction:
Table 5: A Comprehensive comparison of various PDF parsers for table detection across different document categories. For Camelot, pdfplumber, PyMuPDF, Tabula, the Jaccard threshold is 0.75. Intersection over Union (IoU) threshold for TATR @ 60 and TATR @ 70 is 0.60 and 0.70 respectively. The best scores for each category are highlighted in bold. Higher values (indicated by ↑) are better for all metrics. The evaluation used 400 balanced documents per category, ensuring a fair comparison across different document types.
@@ 5 Discussion
For text extraction, all rule-based parsers underperformed in the Scientific and Patent categories. Scientific documents are challenging to parse due to the Mathematical expressions in them. All of the rule-based parsers analyzed here extract the mathematical equation in symbolic form. We found that it is insufficient to express the complex formulas involving vectors, matrices, etc. Also, the parsers sometimes extract the content of graphs and it gets mixed in paragraphs which further worsens the quality of extraction. In our opinion, the better way would be to use parsers that can extract the texts from scientific documents as Latex, Markdown, or MathML formats. As we show with an example of Nougat, these approaches are much better for scientific documents than the conventional rule-based approaches. In the DocLayNet dataset, most patents are documents filled with images of designs or diagrams of chemical compounds, etc. Parsing such documents is beyond the scope of rule-based parsers. Hence OCR based approach would be more suitable for such documents.
Table detection is a challenging part for all of the parsers. Rule-based parsers excel in detecting the table if i) there are clear boundaries in tables and ii) Spaces between columns’ text are fixed. However, most of the parsers in this category were not able to parse the PDF at all. It can be seen in low recall/high false positive scores. However, the diversity in table structures across document categories poses a considerable obstacle for rule-based parsers. We found that all of these parsers mainly struggle to detect the tables if i) Tables are nested ii) There is no table boundary iii) the Table is in the form of the table of contents iv) Multiple tables on a single page v) Tables using color differentiation instead of lines vi) Table columns are separated by “..” or “ - -”. We found that the learning-based approach would be more suitable for handling complex tables. As can be seen in Table 5 - TATR although trained for financial and scientific documents also excels in other categories.
@@ 6 Conclusion and Future work
While our study provides valuable insights into the performance of various PDF parsing tools across different document categories it has a few limitations - The sample size for comparison of Nougat is relatively small due to the nature of the ground truth of the DocLayNet dataset(Ground truth is not in LaTex format). One major challenge was dealing with scientific documents, which often contain inline equations which are Mathematical expressions embedded within the text and crucial for the document’s integrity. However, the dataset we used didn’t support extracting inline equations, making accurate text comparison difficult. Also, comparison to learning-based methods was limited to a single tool each for scientific text extraction (Nougat) and table detection (TATR).

### v2

@@ 2 Related Work
We first survey the existing labeled datasets for information extraction tasks from PDFs. The earliest datasets for Document Image Analysis and Recognition (DIAR) can be traced back to the 1990s, including NIST[14] and UW[15] datasets. Over the past 30 years, there has been significant progress in this direction. Notably, since 2015, the field has seen a boom in terms of dataset availability. Initially, the datasets were small and comprised scanned PDFs or images, but as time progressed, datasets with digitally born PDFs became more prevalent. In creating a Document Layout Analysis (DLA) dataset, annotation is the most challenging part. The process of annotation can be broadly classified into three categories[16]:
@@
Table 2: Summary of studies comparing PDF parsers. Evaluation metrics used by the studies: Precision(P), Recall(R), and F1 Score.
@@ 4.1 For Text Extraction:
Scientific and Patent categories presented notable challenges. In the Scientific category, all tools showed a marked decrease in performance, with Pypdfium maintaining a slight edge at an F1 score of 0.8525 and BLEU score of 0.7089, as shown in Table 4. The Patent category exhibited the widest performance gap among tools. PyMuPDF and pypdfium significantly outperformed others, scoring F1 scores of 0.973 and 0.969 respectively. The significant drop in scores for Scientific and Patent documents highlights a persistent challenge in PDF parsing technology. For Scientific documents, we compare the rule-based parser with Nougat a Visual transformer model and it outperforms all rule-based parsers by a huge margin as shown in Figure 8.
@@ 4.2 For Table Extraction:
Table 5: A comprehensive comparison of various PDF parsers for table detection across different document categories. For Camelot, pdfplumber, PyMuPDF, and Tabula, the Jaccard threshold is 0.75. Intersection over Union (IoU) threshold for TATR @ 60 and TATR @ 70 is 0.60 and 0.70, respectively. The best scores for each category are highlighted in bold. Higher values (indicated by ↑) are better for all metrics. The evaluation used 400 balanced documents per category, ensuring a fair comparison across different document types.
@@ 5 Discussion
For text extraction, all rule-based parsers underperformed in the Scientific and Patent categories. Scientific documents are challenging to parse due to the Mathematical expressions in them. All of the rule-based parsers analyzed here extract the mathematical equation in symbolic form. We found that it is insufficient to express the complex formulas involving vectors, matrices, etc. Also, the parsers sometimes extract the content of graphs and it gets mixed in paragraphs which further worsens the quality of extraction. In our opinion, the better way would be to use parsers that can extract the texts from scientific documents as Latex, Markdown, or MathML formats. As we show with an example of Nougat, these approaches are much better for scientific documents than the conventional rule-based approaches. In the DocLayNet dataset, most patents are documents filled with images of designs or diagrams of chemical compounds, etc. Parsing such documents is beyond the scope of rule-based parsers. Hence, an OCR-based approach would be more suitable for such documents.
Table detection is a challenging part for all of the parsers. Rule-based parsers excel in detecting the table if i) there are clear boundaries in tables and ii) Spaces between columns’ text are fixed. However, most of the parsers in this category were not able to parse the PDF at all. It can be seen in low recall/high false positive scores. However, the diversity in table structures across document categories poses a considerable obstacle for rule-based parsers. We found that all of these parsers mainly struggle to detect the tables if i) Tables are nested ii) There is no table boundary iii) the Table is in the form of the table of contents iv) Multiple tables on a single page v) Tables using color differentiation instead of lines vi) Table columns are separated by “..” or “ - -”. We found that the learning-based approach would be more suitable for handling complex tables. As can be seen in Table 5 - TATR although trained for financial and scientific documents, also excels in other categories.
@@ 6 Conclusion and Future Work
While our study provides valuable insights into the performance of various PDF parsing tools across different document categories, it has a few limitations - The sample size for comparison of Nougat is relatively small due to the nature of the ground truth of the DocLayNet dataset(Ground truth is not in LaTex format). One major challenge was dealing with scientific documents, which often contain inline equations, which are Mathematical expressions embedded within the text and crucial for the document’s integrity. However, the dataset we used didn’t support extracting inline equations, making accurate text comparison difficult. Also, comparison to learning-based methods was limited to a single tool each for scientific text extraction (Nougat) and table detection (TATR).